An important application of regression analysis in accounting is in the estimation of cost. By collecting data on volume and cost and using the least squares method to develop an estimated regression equation relating volume and cost, an accountant can estimate the cost associated with a particular manufacturing volume. Consider the following sample of production volumes and total cost data for a manufacturing operation.

Production Volume (units) Total Cost ($)
400 3800
500 4300
600 5300
650 5900
750 6300
800 7100

The data on the production volume  and total cost  for particular manufacturing operation were used to develop the estimated regression equation.

a. The company’s production schedule shows that  units must be produced next month. What is the point estimate of the total cost for next month?

  (to 2 decimals)

b. Develop a  prediction interval for the total cost for next month.

  (to 2 decimals)
-value   (to 3 decimals)
  (to 2 decimals)

Prediction Interval for an Individual Value:

( ,   ) (to whole number)An important application of regression analysis in accounting is in the estimation of cost. By collecting data on volume and
Answer #1

a)

Answer:

4905.88

Explanation:

The regression analysis is done in excel by following these steps,

Step 1: Write the data values in excel,

Step 2: DATA > Data Analysis > Regression > OK. The screenshot is shown below,

IE INSERT PAGE LAYOUT FORMULAS DATA REVIEW 000 - Data Analysis Х ? OK Cancel AZA Data Analysis Analysis Tools Fourier Analysi

Step 3: Select Input Y Range: 'Y' column, Input X Range: 'X' column then OK. The screenshot is shown below,

A1 X ✓ fx A B C E F H — 1 x iy 4001 2 ! 3 ? X 4 500! 600! 650 750 OK 5 $B$1:$B$7 S 3800 4300 Regression 5300 Input 5900 Input

The result is obtained. The screenshot is shown below,

J K L M N 0 Р 1 SUMMARY OUTPUT 2 3 Regression Statistics 4 Multiple R 0.989225 5 R Square 0.978565 6 Adjusted R Square 0.9732

The regression equation is,

\widehat{Y}=416.9118+8.1618X

For X = 550

\widehat{Y}=416.9118+8.1618\times 550=4905.882

d)

Answer:

s=203.33

t-\text{value}=2.776

s_{\text{Pred}} =223.28

\text{95\% PI}=(4285.96,5525.80)

Explanation:

From the regression output summary,

The standard error of the regression is,

s=203.3271\approx 203.33

The t-critical value is obtained from t-distribution table for significance level = 0.05 and degree of freedom = n - 2 = 6 - 2 = 4

t_{1-\alpha/2}=2.7764

The standard error of the prediction is obtained using the following formula,

s_{\text{Pred}} =s\sqrt{1+\frac{1}{n}+\frac{(X_{\text{predictor}}-\overline{X})^2}{\sum X^2-\frac{1}{n}(\sum X)^2}}

From the data values,

x x^2
400 160000
500 250000
600 360000
650 422500
750 562500
800 640000
Sum=3700 2395000

\overline{X}=\frac{3700}{6}=616.6667

\sum X^2=2395000

s_{\text{Pred}} =s\sqrt{1+\frac{1}{6}+\frac{(550-616.6667)^2}{2395000-\frac{1}{6} (3700)^2}}

s_{\text{Pred}} =223.2789

Now, the confidence interval for Xpredictor = 550 is obtained using the formula,

\text{PI}=\widehat{Y}\pm t_{1-\alpha/2}\times s_{\text{Pred}}

\text{PI}=4905.882\pm 2.7764\times 223.2789

\text{95\% PI}=(4285.96,5525.80)